OpenStudy (anonymous):

12P2 evaluate the expression

OpenStudy (anonymous):

\[Evaluate the expression 12P2\]

OpenStudy (anonymous):

Do you mean \(12p^2\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

There is nothing to evaluate there.

OpenStudy (anonymous):

OK! Thank You gypsy1274

OpenStudy (anonymous):

\[ ^{12}P_2=\frac {12!}{(12-2)!}=\frac{12!}{10!}=\frac{12\times 11\times 10!}{10!}=12\times11=132 \]

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."